# Parameter-Efficient and Personalized Federated Training of Generative Models at the Edge

### v1

@@ 1 Introduction
Large generative models enable high-fidelity text, vision, and multimodal generation but pose privacy risks because gradients can leak training data zhu2019deep . Differential privacy provides formal protection for iterative training mcmahan2018differentially . Federated learning (FL) trains collaboratively without centralizing raw data mcmahan2017communication . Surveys synthesize advances and open problems across systems and theory kairouz2021advances . A practical framework eases FL experimentation at scale beutel2020flower . Wireless and edge computing highlight deployment constraints and opportunities niknam2020federated . Multi-institutional medicine shows the feasibility of privacy-preserving collaboration sheller2020federated .
@@
System heterogeneity motivates proximal objectives that stabilize partial or variable client work li2020federated . Control variates correct client drift under non-IID data karimireddy2020scaffold . Objective normalization addresses unequal local steps wang2020tackling . Hierarchical aggregation reduces wall-clock time on tiered edges wang2021resource . Guided participant selection improves time-to-accuracy via utility-aware sampling lai2021oort . Experience-driven model migration accelerates convergence in heterogeneous networks 10508334 . Asynchronous protocols reduce idle waiting under bandwidth variance liu2021communication . Probabilistic communication improves decentralized throughput without central coordination liu2025accelerating . Adaptive asynchronous training addresses resource-constrained clients liu2021adaptive . Progressive semi-supervised curricula enhance label efficiency on devices liu2024enhancing . Layer-wise aggregation supports partially connected or decentralized topologies liu2023yoga . Block-wise regularization with distillation improves robustness under heterogeneity liu2023adaptive . Neural composition with adaptive local updates further accelerates convergence liu2025adaptive .
@@
For large models, parameter-efficient fine-tuning (PEFT) uses lightweight adapters to reduce trainable parameters houlsby2019parameter . Prompt tuning scales efficient adaptation to large backbones lester2021power . Prefix-tuning optimizes continuous prompts for generation li2021prefix . LoRA forms compact low-rank updates for dense layers hu2021lora . FedLoRA adapts LoRA to the federated setting zhang2023fedlora . FedPrompt communicates prompts efficiently under privacy constraints lu2023fedprompt . FedTuning provides a general federated PEFT framework for LLMs zhang2024fedtuning . Data-free distillation mitigates non-IID without sharing examples zhu2021data . Ensemble distillation improves model fusion robustness lin2020ensemble . Mobile diffusion exposes strict latency/energy budgets for on-device generation lee2024snapfusion . Robust token sampling (Top-$n\sigma$) stabilizes decoding under noise tang2025topnSigma . Federated fine-tuning benefits from adaptive LoRA placement with activation quantization li2025fedquad . End–cloud pipeline optimization speeds collaborative inference gao2025accelerating . Expert-split scheduling accelerates MoE inference on the edge yan2025accelerating . Federated LLM systems engineering continues to mature fan2024fedllm . Early FL with pre-trained LMs laid a foundation for later systems lin2021fedbert . Privacy-preserving federated diffusion extends generative FL beyond language xiong2023feddiffusion . Personalized diffusion tailors generative models to clients’ data shoshan2024pfd . GAN-based federation provides a complementary generative path rasouli2020fedgan .
@@
Security remains central because backdoor/poisoning can corrupt the global model bagdasaryan2020how . Fairness-aware or agnostic risk minimization improves worst-group robustness mohri2019agnostic . Split learning trades different privacy–compute envelopes than FL vepakomma2018split . Federated unlearning supports the “right to be forgotten” after training wu2022federated . Federated reinforcement learning extends collaboration to control tasks zhu2021federated . Attention-based federation benefits multi-agent coordination zhang2021fedformer . Federated unsupervised representation learning reduces dependence on labeled data van2022federated . Federated graph learning targets structural non-IID challenges wu2022fedgraph .
@@
Privacy-sensitive edge sensing motivates our setting. Anti-interference WiFi HAR began with subcarrier-correlation selection 9075376 . Interference-independent phase modeling improved robustness to co-channel noise 9613773 . Attention-based WiFi gestures enabled fine-grained HCI on commodity devices 10090421 . Multimodal WiFi-vision emotion recognition broadened affective computing 10149418 . Open-set WiFi gestures addressed deployment-time uncertainty 10899398 . Commodity WiFi measured pulmonary function without mouth clinging 10699348 . Target-oriented in-area sensing advanced respiratory healthcare scenarios 10804189 . RFID sensing recognized handwriting and identity ubiquitously feng2025rf . Smartphone acoustics demonstrated keystroke eavesdropping risks 10587029 . PHY-layer fingerprint-based authentication faced stealthy attacks that bypass fingerprints 10339891 . Distribution shift in dynamic FER called for robust optimization cui2025learningheterogeneitygeneralizingdynamic . Label-noise suppression improved FER reliability under noisy supervision 10916982 . Efficient temporal filtering strengthened video grounding with dynamics hu2025unified . Communication-efficient decentralized graph FL addressed non-IID structures at scale wang2025towards .
@@ 2.1 Foundations, Heterogeneity, and Optimization
FL enables training on decentralized data without raw-data centralization mcmahan2017communication . A comprehensive survey consolidates open problems and progress kairouz2021advances . Wireless/edge perspectives highlight networked constraints niknam2020federated . In healthcare, FL enables cross-site modeling under privacy regulations sheller2020federated . Proximal objectives stabilize updates under heterogeneous devices and partial work li2020federated . Control variates mitigate client drift for non-IID data karimireddy2020scaffold . Objective normalization handles unequal local steps wang2020tackling . Hierarchical aggregation reduces wall-clock time by exploiting edge tiers wang2021resource . Guided participant selection improves time-to-accuracy in practice lai2021oort . Experience-driven model migration accelerates progress on heterogeneous edges 10508334 . Asynchronous protocols reduce idle time in fluctuating bandwidth liu2021communication . Probabilistic communication increases decentralized throughput at scale liu2025accelerating . Adaptive asynchronous training tackles resource-constrained clients liu2021adaptive . Progressive semi-supervised training improves sample efficiency on devices liu2024enhancing . Layer-wise aggregation supports decentralized or partially connected overlays liu2023yoga . Block-wise regularization with distillation enhances robustness under heterogeneity liu2023adaptive . Neural composition and adaptive local updates further accelerate convergence liu2025adaptive .
@@ 2.2 Personalization and Architecture Adaptation
A simple but effective personalization strategy partitions models into shared backbones and private heads arivazhagan2019federated . Meta-learning adapts rapidly with guarantees in federated settings fallah2020personalized . Proximal personalization balances local and global optima li2021ditto . Ferrari targets robust personalization for heterogeneous edge clients yao2024ferrari . NAS-driven personalization explores client-aware designs at the edge yan2024peaches . Heterogeneity-aware federated NAS discovers architectures tailored to client clusters 10251628 . Earlier federated NAS established search under FL constraints he2020fednas . Federated graph learning copes with structural non-IID data distributions wu2022fedgraph . Federated reinforcement learning extends beyond supervised learning to control problems zhu2021federated . Attention-based federation benefits multi-agent coordination zhang2021fedformer . Federated unsupervised representation learning mitigates label scarcity van2022federated .
@@ 2.3 PEFT for Large Models and Federated Generative Learning
Adapter-based PEFT reduces trainable parameters dramatically houlsby2019parameter . Prompt tuning scales parameter-efficient adaptation to large backbones lester2021power . Prefix-tuning optimizes continuous prompts for generation li2021prefix . LoRA provides compact low-rank updates for dense layers hu2021lora . FedLoRA transfers LoRA’s benefits to federated training zhang2023fedlora . FedPrompt communicates compact prompts under privacy constraints lu2023fedprompt . FedTuning systematizes PEFT choices for federated LLMs zhang2024fedtuning . Data-free distillation mitigates non-IID without raw-data sharing zhu2021data . Ensemble distillation improves robust fusion of client models lin2020ensemble . On-device diffusion reveals tight latency and energy limits on modern phones lee2024snapfusion . Robust token sampling (Top-$n\sigma$) stabilizes decoding in noisy regimes tang2025topnSigma . Adaptive LoRA placement plus activation quantization cut communication during federated fine-tuning li2025fedquad . End–cloud pipeline optimization boosts collaborative inference throughput gao2025accelerating . Expert-split scheduling accelerates MoE inference at the network edge yan2025accelerating . Federated LLM systems continue to mature with systems-level engineering fan2024fedllm . Early work federating pre-trained LMs paved the way for later large-scale systems lin2021fedbert . Privacy-preserving federated diffusion broadens beyond language xiong2023feddiffusion . Personalized diffusion tailors generative models for specific clients shoshan2024pfd . GAN-based federation offers a complementary route to diffusion and LLMs rasouli2020fedgan .
@@ 2.4 Security, Privacy, and Unlearning
Gradients can be inverted to reconstruct sensitive examples zhu2019deep . Differential privacy provides statistical guarantees for iterative training mcmahan2018differentially . Backdoor/poisoning attacks threaten model integrity in FL bagdasaryan2020how . Agnostic/fairness-aware FL improves worst-group risk and robustness mohri2019agnostic . Split learning trades different privacy–compute envelopes than FL vepakomma2018split . Federated unlearning enables removal of users’ contributions post-training wu2022federated . Catastrophic forgetting during federated fine-tuning can be mitigated by adaptive transformer block expansion huo2025mitigating .
@@ 2.5 Applications: Radio Sensing and Vision
Anti-interference WiFi HAR started with subcarrier-correlation selection on CSI 9075376 . Interference-independent phase modeling delivers robust activity recognition 9613773 . Attention-based WiFi gestures enable precise HCI on commodity devices 10090421 . WiFi–vision integration supports unobtrusive emotion recognition 10149418 . Open-set WiFi gesture recognition addresses uncertainty in the wild 10899398 . Commodity WiFi can measure pulmonary function without mouthpieces 10699348 . Target-oriented in-area sensing supports respiratory healthcare scenarios 10804189 . RFID sensing recognizes handwriting and identity ubiquitously feng2025rf . Acoustic keystroke eavesdropping exposes side-channel risks on smartphones 10587029 . PHY-layer fingerprint-based authentication can be bypassed by undetectable attacks 10339891 . Robust dynamic FER improves cross-domain generalization cui2025learningheterogeneitygeneralizingdynamic . Label-noise suppression improves FER reliability with noisy supervision 10916982 . Efficient temporal filtering benefits video grounding under motion and distractors hu2025unified . Communication-efficient decentralized graph FL addresses structural non-IID at scale wang2025towards . Federated super-resolution b
@@ 3.2.2 Personalized Adaptation with Low-Rank Adaptation (LoRA)
For the client-specific adaptation modules, we employ Low-Rank Adaptation (LoRA) hu2021lora , a highly effective and efficient PEFT technique. LoRA is based on the hypothesis that the change in weights during model adaptation has a low "intrinsic rank." Instead of directly updating a large weight matrix $W_{0}\in\mathbb{R}^{d\times k}$ from the pre-trained model, LoRA introduces two much smaller, low-rank matrices, $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$, where the rank $r\ll\min(d,k)$. The update to the weight matrix is then represented by the product of these two smaller matrices, $\Delta W=BA$. During fine-tuning, the original weights $W_{0}$ are frozen, and only the parameters of $A$ and $B$ are trained. The forward pass is modified as follows:
@@ 5.1.1 Text Generation on Penn Treebank
The performance on the language modeling task is summarized in Table LABEL:tab:ptb_results. The key metric is Perplexity (PPL) on the global test set, where lower values are better. The Centralized (Full Model) training, which has access to all data in a single location, establishes a strong performance upper bound with a PPL of 75.4. Interestingly, the Centralized (LoRA) baseline, which trains only the lightweight adapter on the same pooled data, achieves a remarkably close PPL of 78.9. This initial result is significant, as it validates our core hypothesis: a small, parameter-efficient module is sufficient to adapt a large pre-trained backbone for a specific domain, echoing the findings of the original LoRA paper hu2021lora .
@@ 5.1.2 Image Generation on CIFAR-10
The results for the conditional image generation task on CIFAR-10, measured by the Fréchet Inception Distance (FID), are presented in Table LABEL:tab:cifar_results. A lower FID score signifies higher-quality and more diverse generated images. Similar to the text task, the Centralized (Full Model) training sets a strong benchmark with an FID of 25.8. The Centralized (LoRA) approach again proves highly effective, achieving a competitive FID of 28.1, confirming that parameter-efficient adaptation is also viable for generative vision models.
@@ 5.2 Analysis of Communication and System Efficiency
A primary motivation for FedGen-Edge is to make the federated training of large models practical for edge environments. We now analyze the communication costs and convergence behavior, which are visualized in Fig. 2 and Fig. LABEL:fig:convergence.
@@
The convergence plots in Fig. LABEL:fig:convergence provide further insights. For the PTB task, we observe that FedGen-Edge not only converges to a much better final perplexity than FedAvg but also does so with a smoother and more stable trajectory. The FedAvg curve exhibits high variance and plateaus early, indicative of the aforementioned client drift problem. FedGen-Edge, however, shows steady improvement over the communication rounds. This stability suggests that our method is more robust to the Non-IID nature of the data. While advanced optimizers like SCAFFOLD karimireddy2020scaffold or FedNova could potentially improve the stability of FedAvg, they do not address the fundamental communication bottleneck. Our approach tackles the communication problem head-on while also inherently regularizing the training, providing a dual benefit. The system-level benefits of reducing communication also align with the goals of frameworks like Oort lai2021oort , as faster uploads from clients would lead to shorter round times and faster overall convergence.

### v3

@@ 1 Introduction
Large generative models enable high-fidelity text, vision, and multimodal generation but pose privacy risks because gradients can leak training data. Differential privacy provides formal protection for iterative training mcmahan2018differentially. Federated learning (FL) trains collaboratively without centralizing raw data mcmahan2017communication. Surveys synthesize advances and open problems across systems and theory kairouz2021advances. A practical framework eases FL experimentation at scale. Wireless and edge computing highlight deployment constraints and opportunities niknam2020federated. Multi-institutional medicine shows the feasibility of privacy-preserving collaboration sheller2020federated.
@@
System heterogeneity motivates proximal objectives that stabilize partial or variable client work li2020federated. Control variates correct client drift under non-IID data karimireddy2020scaffold. Objective normalization addresses unequal local steps wang2020tackling. Hierarchical aggregation reduces wall-clock time on tiered edges wang2021resource. Guided participant selection improves time-to-accuracy via utility-aware sampling lai2021oort. Experience-driven model migration accelerates convergence in heterogeneous networks. Asynchronous protocols reduce idle waiting under bandwidth variance liu2021communication. Probabilistic communication improves decentralized throughput without central coordination. Adaptive asynchronous training addresses resource-constrained clients. Progressive semi-supervised curricula enhance label efficiency on devices liu2024enhancing. Layer-wise aggregation supports partially connected or decentralized topologies liu2023yoga. Block-wise regularization with distillation improves robustness under heterogeneity liu2023adaptive. Neural composition with adaptive local updates further accelerates convergence liu2025adaptive.
@@
For large models, parameter-efficient fine-tuning (PEFT) uses lightweight adapters to reduce trainable parameters houlsby2019parameter. Prompt tuning scales efficient adaptation to large backbones lester2021power. Prefix-tuning optimizes continuous prompts for generation li2021prefix. LoRA forms compact low-rank updates for dense layers. FedLoRA adapts LoRA to the federated setting zhang2023fedlora. FedPrompt communicates prompts efficiently under privacy constraints lu2023fedprompt. FedTuning provides a general federated PEFT framework for LLMs zhang2024fedtuning. Data-free distillation mitigates non-IID without sharing examples zhu2021data. Ensemble distillation improves model fusion robustness lin2020ensemble. Mobile diffusion exposes strict latency/energy budgets for on-device generation. Robust token sampling (Top-$n\sigma$) stabilizes decoding under noise tang2025topnSigma. Federated fine-tuning benefits from adaptive LoRA placement with activation quantization li2025fedquad. End–cloud pipeline optimization speeds collaborative inference gao2025accelerating. Expert-split scheduling accelerates MoE inference on the edge yan2025accelerating. Federated LLM systems engineering continues to mature fan2024fedllm. Early FL with pre-trained LMs laid a foundation for later systems lin2021fedbert. Privacy-preserving federated diffusion extends generative FL beyond language xiong2023feddiffusion. Personalized diffusion tailors generative models to clients’ data shoshan2024pfd. GAN-based federation provides a complementary generative path rasouli2020fedgan.
@@
Security remains central because backdoor/poisoning can corrupt the global model bagdasaryan2020how. Fairness-aware or agnostic risk minimization improves worst-group robustness mohri2019agnostic. Split learning trades different privacy–compute envelopes than FL. Federated unlearning supports the “right to be forgotten” after training wu2022federated. Federated reinforcement learning extends collaboration to control tasks zhu2021federated. Attention-based federation benefits multi-agent coordination zhang2021fedformer. Federated unsupervised representation learning reduces dependence on labeled data. Federated graph learning targets structural non-IID challenges wu2022fedgraph.
@@
Privacy-sensitive edge sensing motivates our setting. Anti-interference WiFi HAR began with subcarrier-correlation selection. Interference-independent phase modeling improved robustness to co-channel noise. Attention-based WiFi gestures enabled fine-grained HCI on commodity devices Multimodal WiFi-vision emotion recognition broadened affective computing 10149418. Open-set WiFi gestures addressed deployment-time uncertainty. Commodity WiFi measured pulmonary function without mouth clinging. Target-oriented in-area sensing advanced respiratory healthcare scenarios. RFID sensing recognized handwriting and identity ubiquitously. Smartphone acoustics demonstrated keystroke eavesdropping risks. PHY-layer fingerprint-based authentication faced stealthy attacks that bypass fingerprints. Distribution shift in dynamic FER called for robust optimization. Label-noise suppression improved FER reliability under noisy supervision. Efficient temporal filtering strengthened video grounding with dynamics. Communication-efficient decentralized graph FL addressed non-IID structures at scale wang2025towards.
@@ 2.1 Foundations, Heterogeneity, and Optimization
FL enables training on decentralized data without raw-data centralization mcmahan2017communication. A comprehensive survey consolidates open problems and progress kairouz2021advances. Wireless/edge perspectives highlight networked constraints niknam2020federated. In healthcare, FL enables cross-site modeling under privacy regulations sheller2020federated. Proximal objectives stabilize updates under heterogeneous devices and partial work li2020federated. Control variates mitigate client drift for non-IID data karimireddy2020scaffold. Objective normalization handles unequal local steps wang2020tackling. Hierarchical aggregation reduces wall-clock time by exploiting edge tiers wang2021resource. Guided participant selection improves time-to-accuracy in practice lai2021oort. Experience-driven model migration accelerates progress on heterogeneous edges. Asynchronous protocols reduce idle time in fluctuating bandwidth liu2021communication. Probabilistic communication increases decentralized throughput at scale. Adaptive asynchronous training tackles resource-constrained clients. Progressive semi-supervised training improves sample efficiency on devices liu2024enhancing. Layer-wise aggregation supports decentralized or partially connected overlays liu2023yoga. Block-wise regularization with distillation enhances robustness under heterogeneity liu2023adaptive. Neural composition and adaptive local updates further accelerate convergence liu2025adaptive.
@@ 2.2 Personalization and Architecture Adaptation
A simple but effective personalization strategy partitions models into shared backbones and private heads arivazhagan2019federated. Meta-learning adapts rapidly with guarantees in federated settings fallah2020personalized. Proximal personalization balances local and global optima li2021ditto. Ferrari targets robust personalization for heterogeneous edge clients yao2024ferrari. NAS-driven personalization explores client-aware designs at the edge yan2024peaches. Heterogeneity-aware federated NAS discovers architectures tailored to client clusters. Earlier federated NAS established search under FL constraints he2020fednas. Federated graph learning copes with structural non-IID data distributions wu2022fedgraph. Federated reinforcement learning extends beyond supervised learning to control problems zhu2021federated. Attention-based federation benefits multi-agent coordination zhang2021fedformer. Federated unsupervised representation learning mitigates label scarcity.
@@ 2.3 PEFT for Large Models and Federated Generative Learning
Adapter-based PEFT reduces trainable parameters dramatically houlsby2019parameter. Prompt tuning scales parameter-efficient adaptation to large backbones lester2021power. Prefix-tuning optimizes continuous prompts for generation li2021prefix. LoRA provides compact low-rank updates for dense layers. FedLoRA transfers LoRA’s benefits to federated training zhang2023fedlora. FedPrompt communicates compact prompts under privacy constraints lu2023fedprompt. FedTuning systematizes PEFT choices for federated LLMs zhang2024fedtuning. Data-free distillation mitigates non-IID without raw-data sharing zhu2021data. Ensemble distillation improves robust fusion of client models lin2020ensemble. On-device diffusion reveals tight latency and energy limits on modern phones. Robust token sampling (Top-$n\sigma$) stabilizes decoding in noisy regimes tang2025topnSigma. Adaptive LoRA placement plus activation quantization cut communication during federated fine-tuning li2025fedquad. End–cloud pipeline optimization boosts collaborative inference throughput gao2025accelerating. Expert-split scheduling accelerates MoE inference at the network edge yan2025accelerating. Federated LLM systems continue to mature with systems-level engineering fan2024fedllm. Early work federating pre-trained LMs paved the way for later large-scale systems lin2021fedbert. Privacy-preserving federated diffusion broadens beyond language xiong2023feddiffusion. Personalized diffusion tailors generative models for specific clients shoshan2024pfd. GAN-based federation offers a complementary route to diffusion and LLMs rasouli2020fedgan.
@@ 2.4 Security, Privacy, and Unlearning
Gradients can be inverted to reconstruct sensitive examples. Differential privacy provides statistical guarantees for iterative training mcmahan2018differentially. Backdoor/poisoning attacks threaten model integrity in FL bagdasaryan2020how. Agnostic/fairness-aware FL improves worst-group risk and robustness mohri2019agnostic. Split learning trades different privacy–compute envelopes than FL. Federated unlearning enables removal of users’ contributions post-training wu2022federated. Catastrophic forgetting during federated fine-tuning can be mitigated by adaptive transformer block expansion huo2025mitigating.
@@ 2.5 Applications: Radio Sensing and Vision
Anti-interference WiFi HAR started with subcarrier-correlation selection on CSI. Interference-independent phase modeling delivers robust activity recognition. Attention-based WiFi gestures enable precise HCI on commodity devices WiFi–vision integration supports unobtrusive emotion recognition. Open-set WiFi gesture recognition addresses uncertainty in the wild. Commodity WiFi can measure pulmonary function without mouthpieces. Target-oriented in-area sensing supports respiratory healthcare scenarios. RFID sensing recognizes handwriting and identity ubiquitously. Acoustic keystroke eavesdropping exposes side-channel risks on smartphones. PHY-layer fingerprint-based authentication can be bypassed by undetectable attacks. Robust dynamic FER improves cross-domain generalization. Label-noise suppression improves FER reliability with noisy supervision. Efficient temporal filtering benefits video grounding under motion and distractors. Communication-efficient decentralized graph FL addresses structural non-IID at scale wang2025towards. Federated super-resolution b
@@ 3.2.2 Personalized Adaptation with Low-Rank Adaptation (LoRA)
For the client-specific adaptation modules, we employ Low-Rank Adaptation (LoRA), a highly effective and efficient PEFT technique. LoRA is based on the hypothesis that the change in weights during model adaptation has a low "intrinsic rank." Instead of directly updating a large weight matrix $W_{0}\in\mathbb{R}^{d\times k}$ from the pre-trained model, LoRA introduces two much smaller, low-rank matrices, $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$, where the rank $r\ll\min(d,k)$. The update to the weight matrix is then represented by the product of these two smaller matrices, $\Delta W=BA$. During fine-tuning, the original weights $W_{0}$ are frozen, and only the parameters of $A$ and $B$ are trained. The forward pass is modified as follows:
@@ 5.1.1 Text Generation on Penn Treebank
The performance on the language modeling task is summarized in Table LABEL:tab:ptb_results. The key metric is Perplexity (PPL) on the global test set, where lower values are better. The Centralized (Full Model) training, which has access to all data in a single location, establishes a strong performance upper bound with a PPL of 75.4. Interestingly, the Centralized (LoRA) baseline, which trains only the lightweight adapter on the same pooled data, achieves a remarkably close PPL of 78.9. This initial result is significant, as it validates our core hypothesis: a small, parameter-efficient module is sufficient to adapt a large pre-trained backbone for a specific domain, echoing the findings of the original LoRA paper.
@@ 5.1.2 Image Generation on CIFAR-10
The results for the conditional image generation task on CIFAR-10, measured by the Fréchet Inception Distance (FID), are presented in Table. A lower FID score signifies higher-quality and more diverse generated images. Similar to the text task, the Centralized (Full Model) training sets a strong benchmark with an FID of 25.8. The Centralized (LoRA) approach again proves highly effective, achieving a competitive FID of 28.1, confirming that parameter-efficient adaptation is also viable for generative vision models.
@@ 5.2 Analysis of Communication and System Efficiency
A primary motivation for FedGen-Edge is to make the federated training of large models practical for edge environments. We now analyze the communication costs and convergence behavior, which are visualized in Fig. 2 and Fig.
@@
The convergence plots in Fig. provide further insights. For the PTB task, we observe that FedGen-Edge not only converges to a much better final perplexity than FedAvg but also does so with a smoother and more stable trajectory. The FedAvg curve exhibits high variance and plateaus early, indicative of the aforementioned client drift problem. FedGen-Edge, however, shows steady improvement over the communication rounds. This stability suggests that our method is more robust to the Non-IID nature of the data. While advanced optimizers like SCAFFOLD karimireddy2020scaffold or FedNova could potentially improve the stability of FedAvg, they do not address the fundamental communication bottleneck. Our approach tackles the communication problem head-on while also inherently regularizing the training, providing a dual benefit. The system-level benefits of reducing communication also align with the goals of frameworks like Oort lai2021oort, as faster uploads from clients would lead to shorter round times and faster overall convergence.